# Quantum Optimization for the Maximum Cut Problem on a Superconducting Quantum Computer

### v1

@@ Acknowledgements.
Table S3: Number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ defined in Eq. (S14), summing all problem sizes and problem instances used in this work, for $p=1$, $p=2$, and $p=3$ QAOA layers. The total number of entries is the sum of $N(N-1)/2\times\textrm{number of instances}$, for each problem size $N$ considered (1000 problem instances were used for all problem sizes, and the sum over $n=5,...,12$ captures contributions from all problem sizes: $40962^{5}=32,...,2^{12}=$ 4096. The total number of nonzero entries is obtained by eliminating all trivially zero entries which arise from nonintersecting light cones. The total number of unique entries accounts for isomorphism between the subgraphs made of intersecting light cones: this is the minimum number of QAOA runs required and the number used in this work.
@@
We show in Tab. S3 the total number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ (defined in Eq. (S14)), summed across all problem instances and sizes $N$ considered in this work. Using the fact that nonintersecting light cones yield a trivially zero correlation between variables and leveraging isomorphism between the subgraphs resulting from intersecting light cones considerably reduces the number of entries. For instance, for $p=1$, one only needs to run $44$ different QAOA circuits to build and fill the correlation matrices of all the $8,000$ randomly generated problem instances considered in this work, from size $N=32$ to $N=4,096$. This greatly facilitates the collection of statistics required to evaluate the performance of quantum solvers over many problem instances and problem sizes. Moving to $p=2$, one needs to execute $70,425$ circuits to solve $8,000$ randomly generated problem instances from size $N=32$ to $N=4,096$.
Average number of entries versus N
@@
Figure S6: (a) Average number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ defined in Eq. (S14) for $p=1$, $p=2$, and $p=3$ QAOA layers, as a function of the problem size $N$. There are naively $O(N^{2})$ entries. Based on the light-cone picture from the QAOA, there are $O(N)$ nonzero entries. Using graph isomorphism, there are $O(1)$ unique entries. (b) Local exponent of the $N$-scaling of the curves in (a), computed using finite differences. The results show the big $O$ scaling with $N$. Error bars indicate one standard deviation.
@@
We show in Fig. S6a the average number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ for a typical problem instance, for $p=1$, $p=2$, and $p=3$ QAOA layers, as a function of the problem size $N$. For a given problem size, there are naively $N(N-1)/2$ entries in the correlation matrix, as it is symmetric.
@@
By only considering entries $(i,j)$ where the light cones centered around $i$ and $j$ intersect, one can reduce this $O(N^{2})$ scaling to $O(N)$, as shown in Fig. S6a. Indeed, in the limit of large $N$, one expects only $O(N)$ nonzero entries when $p\ll\ln N$, because the average shortest path between two vertices grows as $\ln N$ for $3$-regular graphs (see Fig. S2). Each of these nonzero entries corresponds to a subgraph arising from the intersection of the light cones centered around vertices $i$ and $j$.
@@
Among these $O(N)$ subgraphs, many of them turn out to be identical in the large $N$ limit for $p\ll\ln N$. We observe in Fig. S6a that only a constant fraction of the subgraphs are different in a typical problem instance of size $N$. Furthermore, the number of unique subgraphs is independent of the problem size. For instance, for $p=1$ QAOA layer, only about $8$ entries in the correlation matrix are on average nonzero and unique. Hence, in about $8$ QAOA runs one can build the full $N\times N$ correlation matrix for a typical problem instance. The big $O$ scaling is valid in the limit of $N\to+\infty$. For completeness, we plot in Fig. S6b the $N$-dependent local exponent $\alpha(N)$ of $O(N{{}^{\alpha}(N)})$.
@@
The aforementioned scaling can be understood as follows. In the large $N$ limit: (i) the majority of the subgraphs saturate the maximum size bound of Eq. (S20) (Fig. S4); and (ii) the majority of the subgraphs have a treewidth of one (Fig. S5), which shows that the majority of the subgraphs are trees. Combining both observations it becomes clear that most of the subgraphs consist of two $3$-Cayley trees connected by a single common vertex. An example is shown in see Fig. S3(e). There is only one such graph for a given number of QAOA layers $p$. Therefore the majority of the subgraphs, in the large $N$ limit, are isomorphic.
S3.3.2 Determining (Sub)Graph Isomorphism
@@
    $\bigwedge\nolimits_{i=1}^{|V|}\Bigl{(}\bigvee\nolimits_{j=1}^{|V^{\prime}|}x_{i,j}\Bigr{)}~{}\land~{}\bigwedge\nolimits_{i=1}^{|V|}\Bigl{(}\bigvee\nolimits_{j=2,k=1,j>k}^{|V^{\prime}|}\neg x_{i,j}\lor\neg x_{i,k}\Bigr{)}.$ (S21)
@@
    $\bigwedge\nolimits_{k=1}^{|V^{\prime}|}\Bigl{(}\bigvee\nolimits_{j=1,i=2,i>j}^{|V|}\neg x_{i,k}\lor\neg x_{j,k}\Bigr{)}.$ (S22)
@@
    $\bigwedge\nolimits_{(i,j)\in E,k=1}^{|V^{\prime}|}\Bigl{[}\neg x_{i,k}\lor\Bigl{(}\bigvee\nolimits_{(k,k^{\prime})\in E^{\prime}}x_{j,k^{\prime}}\Bigr{)}\Bigr{]}.$ (S23)
The conjunction of the three conditions from Eqs. (S21), (S22), and  (S23) makes a SAT formulation of the graph isomorphism problem. The more general case with $|V|\geq|V^{\prime}|$ is known as the subgraph isomorphism problem, asking whether $G$ contains a subgraph that is isomorphic to $G^{\prime}$. The subgraph isomorphism problem is known to be NP-complete.
Practical Implementation
@@
We consider a few randomly generated problems with $N=4,096$ variables to estimate the runtime for the SAT solver to determine whether two subgraphs are isomorphic. The runtime includes building the SAT problem from two input subgraphs and running the SAT solver itself. Following Eq. (S20), the maximum size of a subgraph as a function of the number of QAOA layers $p$ is $1+6(2^{p}-1)$. The distribution of subgraph sizes was discussed in Sec. S3.2.2 and plotted in Fig. S4. Here, we report the runtime to determine isomorphism on a $64$Gb MacBook Pro with an Apple M1 Max chip at fixed $p$, independent of the size of the subgraphs themselves. The results are shown in Fig. S7a. We observe that the average runtime increases roughly exponentially with $p$, meaning that it is linear in the size of the subgraphs.
Figure S7: Cumulative distribution functions of the runtime for the SAT solver to determine: (a) subgraph isomorphism; and (b) an efficient swap network, for different number of QAOA layers $p$. All tests were conducted on a $64$Gb MacBook Pro with an Apple M1 Max chip.
As shown in Fig. S6, for a given $N$-variables graph problem, there are $O(N)$ subgraph isomorphism problems to solve if one desires to find the $O(1)$ unique entries of the correlation matrix. Given the cost of running the SAT solver $O(N)$ times for a problem of size $N$, it may be preferable to execute $O(N)$ QAOA circuits instead. In the context of this work, we leverage subgraph isomorphism across different problem instances—of the same and different sizes—as well as within problem instances at low $p$, providing a huge benefit, as shown in Tab. S3.
Remarks
For estimating the runtime of the quantum algorithm in Figs. 2 and 3 of the main text and Sec. S4, we assume that we have to execute all of the circuits corresponding to all of the $O(N)$ nonzero entries of the correlation matrix. We do not assume graph isomorphism, given its poor scaling. We only use it as a convenience to get statistics over thousands of problem instances in the context of this work.
S4 Executing Quantum Circuits
S4.1 Quantum Simulations
S4.1.1 Hardware Topology and Swap Network
@@
For the QAOA, each qubit encodes a binary variable of the problem of interest. Therefore we seek the mapping between qubits and variables that will minimize the swap overhead needed to encode the phase separator unitary of Eq. (S13). For an arbitrary mapping between $N$ variables and $N$ qubits, an arbitrary QAOA phase separator unitary can be encoded in a swap network of depth $O(N)$ on, for example, a linear or square hardware topology [33, 94].
Efficient Swap Network Finder
We employ the swap network finder introduced in Ref. [90] and successfully used in Ref. [95]. The hardware topology and phase separator are encoded as graphs $G(V,E)$ and $G^{\prime}(V^{\prime},E^{\prime})$, respectively. Vertices $V^{\prime}$ in the phase separator graph correspond to qubits, and each edge in $E^{\prime}$ corresponds to a term $e^{i\gamma\mathsf{W}_{ij}\hat{Z}_{i}\hat{Z}_{j}/2}\neq\hat{I}$ from Eq. (S13), between two qubits $i$ and $j$. The swap network finder is partly about solving a subgraph isomorphism problem. We perform this task using the approach described in Sec. S3.3, based on a Boolean satisfiability (SAT) formulation of the problem. It works as follows: one iteratively updates the connectivity of the hardware topology graph $G$ by applying a set of swap gates. After the swap gates have been applied, one solves a subgraph isomorphism problem on the updated graph. If $G^{\prime}$ can be embedded onto $G$, then one has found a satisfying mapping of variables onto qubits as well as a satisfying set of swap gates, i.e., a satisfying swap network. If $G^{\prime}$ cannot be embedded onto $G$, one applies another set of swap gates to $G$ and repeats the operation until a satisfying swap network is found. It is the responsibility of the user to define the set of swap gates since there is no optimization of the swap network per se. Hence, one starts with no or a small swap network and increases its depth until a satisfying solution is found.
@@
Figure S8: (a) Graph representing a QAOA phase separator made of $N=6$ qubits. (b) Linear chain topology with $6$ qubits and a single layer of swap gates applied on even edges. (c) Mapping of the QAOA phase separator qubits onto the linear chain topology. The extra effective edges introduced by the layer of swap gates are marked in green. A single layer of swap gates enables the embedding of the graph from (a) onto the linear chain topology.
Data
We map an $N$-variable QAOA phase separator onto an $N$-variable line topology. At each iteration of the swap network builder, we alternatively add a layer of parallel swap gates on even and odd edges of the line topology. This operation is repeated until a satisfying swap network is found. After this step, the quantum circuit encoding the desired QAOA phase separator for a linear qubit topology is fixed. This strategy is applied to an $N=6$-qubit QAOA phase separator in Fig. S8. The swap network builder can find a mapping of the graph nodes onto qubits requiring only a single layer of two-qubit gates. The assignment onto a linear chain is as follows: $[6,1,2,4,5,3]$ with two-qubit operations possible on edges $(6,1)$, $(1,2)$, $(2,4)$, $(4,5)$, and $(5,3)$. From there, a layer of swap gates leads to the following arrangement $[1,6,4,2,3,5]$, enabling two-qubit operations on edges $(1,6)$, $(6,4)$, $(4,2)$, $(2,3)$, $(3,5)$. In particular, edges $(6,4)$ and $(2,3)$ are made available through the swap gates, as shown in Fig. S8, enabling the encoding of the graph onto the linear topology of the qubits. The strategy is guaranteed to succeed with at most $N$ alternating even and odd layers of swap gates [33].
The algorithm for finding an efficient swap network requires a SAT solver for establishing graph isomorphism. The runtimes required to find such swap networks are shown in Fig. S7b. Following Eq. (S20), the maximum size of a subgraph as a function of the number of QAOA layers $p$ is $1+6(2^{p}-1)$.The distribution of subgraph sizes was discussed in Sec. S3.2.2 and plotted in Fig. S4. The runtimes in Fig. S7b were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip at fixed $p$, independent of the size of the subgraphs themselves. We used the SAT solver Glucose [91], based on Minisat [92], through the Python package PySAT [93]. Given that one has to solve a SAT problem, the method becomes prohibitively expensive as $p$ increases. We refer the reader to Ref. [90] for further discussion.
@@
Remarks
For estimating the runtime of the quantum algorithm in Figs. 2 and 3, we assume that the circuit is mapped onto a hardware-native linear chain of qubits with a default, worst-case-scenario swap network yielding a total number of layers of two-qubit gates proportional to $Np$ for $N$ qubits and $p$ QAOA layers [33]. The quantum algorithm depth (and runtime) does not assume the use of an efficient swap network. In the main text and Sec. S4, a default, worst-case-scenario swap network on a linear chain of qubit is assumed.
S4.1.2 Compilation
Figure S9: Construction of algorithmic gates in terms of hardware-native gates on Rigetti’s AnkaaTM-2 superconducting quantum chip. One-qubit gates are defined in Eq. (S24). (a) Implementation of $\texttt{Rx}(\phi)$ for arbitrary angles $\phi$. (b) Implementation of the Hadamard gate. (c) Implementation of the two-qubit $\texttt{Rzz}(\phi)$ gate. (d) Implementation of the two-qubit $\texttt{Rzz}(\phi)\times\texttt{SWAP}$ gate.
The native gates on Rigetti’s AnkaaTM-2 are the one-qubit rotation gates about the $x$ axis $\texttt{Rx}(\phi)=\exp(-i\hat{X}\phi/2)$ where $\phi=\pm\pi/2$, and the two-qubit gate $\texttt{ISWAP}=\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$. One-qubit rotations about the $z$ axis, $\texttt{Rz}(\theta\in\mathbb{R})=\exp(-i\hat{Z}\theta/2)$, are implemented virtually. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. Our parameterized circuits have $\texttt{Rx}(\phi)$, $\texttt{Rzz}(\phi)=\exp[-i\phi\hat{Z}\otimes\hat{Z}/2]$, and $\texttt{Rzz}(\phi)\times\texttt{SWAP}$ gates, which are not native to Rigetti’s AnkaaTM-2. We implement these gates using the compilation in Fig. S9 where

### v2

@@ C.3.1 Connection to the Light Cone Picture from the QAOA
Table 3: Number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ defined in Eq. (17), summing all problem sizes and problem instances used in this work, for $p=1$, $p=2$, and $p=3$ QAOA layers. The total number of entries is the sum of $N(N-1)/2\times\textrm{number of instances}$, for each problem size $N$ considered (1000 problem instances were used for all problem sizes, and the sum over $n=5,...,12$ captures contributions from all problem sizes: $40962^{5}=32,...,2^{12}=$ 4096. The total number of nonzero entries is obtained by eliminating all trivially zero entries which arise from nonintersecting light cones. The total number of unique entries accounts for isomorphism between the subgraphs made of intersecting light cones: this is the minimum number of QAOA runs required and the number used in this work.
@@
We show in Tab. 3 the total number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ (defined in Eq. (17)), summed across all problem instances and sizes $N$ considered in this work. Using the fact that nonintersecting light cones yield a trivially zero correlation between variables and leveraging isomorphism between the subgraphs resulting from intersecting light cones considerably reduces the number of entries. For instance, for $p=1$, one only needs to run $44$ different QAOA circuits to build and fill the correlation matrices of all the $8,000$ randomly generated problem instances considered in this work, from size $N=32$ to $N=4,096$. This greatly facilitates the collection of statistics required to evaluate the performance of quantum solvers over many problem instances and problem sizes. Moving to $p=2$, one needs to execute $70,425$ circuits to solve $8,000$ randomly generated problem instances from size $N=32$ to $N=4,096$.
Average number of entries versus N—
@@
Figure 9: (a) Average number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ defined in Eq. (17) for $p=1$, $p=2$, and $p=3$ QAOA layers, as a function of the problem size $N$. There are naively $O(N^{2})$ entries. Based on the light-cone picture from the QAOA, there are $O(N)$ nonzero entries. Using graph isomorphism, there are $O(1)$ unique entries. (b) Local exponent of the $N$-scaling of the curves in (a), computed using finite differences. The results show the big $O$ scaling with $N$. Error bars indicate one standard deviation.
@@
We show in Fig. 9a the average number of entries in the correlation matrix $\mathsf{Z}^{(p)}$ for a typical problem instance, for $p=1$, $p=2$, and $p=3$ QAOA layers, as a function of the problem size $N$. For a given problem size, there are naively $N(N-1)/2$ entries in the correlation matrix, as it is symmetric.
@@
By only considering entries $(i,j)$ where the light cones centered around $i$ and $j$ intersect, one can reduce this $O(N^{2})$ scaling to $O(N)$, as shown in Fig. 9a. Indeed, in the limit of large $N$, one expects only $O(N)$ nonzero entries when $p\ll\ln N$, because the average shortest path between two vertices grows as $\ln N$ for $3$-regular graphs (see Fig. 5). Each of these nonzero entries corresponds to a subgraph arising from the intersection of the light cones centered around vertices $i$ and $j$.
@@
Among these $O(N)$ subgraphs, many of them turn out to be identical in the large $N$ limit for $p\ll\ln N$. We observe in Fig. 9a that only a constant fraction of the subgraphs are different in a typical problem instance of size $N$. Furthermore, the number of unique subgraphs is independent of the problem size. For instance, for $p=1$ QAOA layer, only about $8$ entries in the correlation matrix are on average nonzero and unique. Hence, in about $8$ QAOA runs one can build the full $N\times N$ correlation matrix for a typical problem instance. The big $O$ scaling is valid in the limit of $N\to+\infty$. For completeness, we plot in Fig. 9b the $N$-dependent local exponent $\alpha(N)$ of $O(N{{}^{\alpha}(N)})$.
@@
The aforementioned scaling can be understood as follows. In the large $N$ limit: (i) the majority of the subgraphs saturate the maximum size bound of Eq. (23) (Fig. 7); and (ii) the majority of the subgraphs have a treewidth of one (Fig. 8), which shows that the majority of the subgraphs are trees. Combining both observations it becomes clear that most of the subgraphs consist of two $3$-Cayley trees connected by a single common vertex. An example is shown in see Fig. 6(e). There is only one such graph for a given number of QAOA layers $p$. Therefore the majority of the subgraphs, in the large $N$ limit, are isomorphic.
C.3.2 Determining (Sub)Graph Isomorphism
@@
    $\bigwedge\nolimits_{i=1}^{|V|}\Bigl{(}\bigvee\nolimits_{j=1}^{|V^{\prime}|}x_{i,j}\Bigr{)}~{}\land~{}\bigwedge\nolimits_{i=1}^{|V|}\Bigl{(}\bigvee\nolimits_{j=2,k=1,j>k}^{|V^{\prime}|}\neg x_{i,j}\lor\neg x_{i,k}\Bigr{)}.$ (24)
@@
    $\bigwedge\nolimits_{k=1}^{|V^{\prime}|}\Bigl{(}\bigvee\nolimits_{j=1,i=2,i>j}^{|V|}\neg x_{i,k}\lor\neg x_{j,k}\Bigr{)}.$ (25)
@@
    $\bigwedge\nolimits_{(i,j)\in E,k=1}^{|V^{\prime}|}\Bigl{[}\neg x_{i,k}\lor\Bigl{(}\bigvee\nolimits_{(k,k^{\prime})\in E^{\prime}}x_{j,k^{\prime}}\Bigr{)}\Bigr{]}.$ (26)
The conjunction of the three conditions from Eqs. (24), (25), and  (26) makes a SAT formulation of the graph isomorphism problem. The more general case with $|V|\geq|V^{\prime}|$ is known as the subgraph isomorphism problem, asking whether $G$ contains a subgraph that is isomorphic to $G^{\prime}$. The subgraph isomorphism problem is known to be NP-complete.
Practical Implementation—
@@
We consider a few randomly generated problems with $N=4,096$ variables to estimate the run-time for the SAT solver to determine whether two subgraphs are isomorphic. The run-time includes building the SAT problem from two input subgraphs and running the SAT solver itself. Following Eq. (23), the maximum size of a subgraph as a function of the number of QAOA layers $p$ is $1+6(2^{p}-1)$. The distribution of subgraph sizes was discussed in App. C.2.2 and plotted in Fig. 7. Here, we report the run-time to determine isomorphism on a $64$Gb MacBook Pro with an Apple M1 Max chip at fixed $p$, independent of the size of the subgraphs themselves. The results are shown in Fig. 10a. We observe that the average run-time increases roughly exponentially with $p$, meaning that it is linear in the size of the subgraphs.
Figure 10: Cumulative distribution functions of the run-time for the SAT solver to determine: (a) subgraph isomorphism; and (b) an efficient swap network, for different number of QAOA layers $p$. All tests were conducted on a $64$Gb MacBook Pro with an Apple M1 Max chip.
As shown in Fig. 9, for a given $N$-variables graph problem, there are $O(N)$ subgraph isomorphism problems to solve if one desires to find the $O(1)$ unique entries of the correlation matrix. Given the cost of running the SAT solver $O(N)$ times for a problem of size $N$, it may be preferable to execute $O(N)$ QAOA circuits instead. In the context of this work, we leverage subgraph isomorphism across different problem instances—of the same and different sizes—as well as within problem instances at low $p$, providing a huge benefit, as shown in Tab. 3.
Remarks—
For estimating the run-time of the quantum algorithm in Figs. 2 and 3 of the main text and App. D, we assume that we have to execute all of the circuits corresponding to all of the $O(N)$ nonzero entries of the correlation matrix. We do not assume graph isomorphism, given its poor scaling. We only use it as a convenience to get statistics over thousands of problem instances in the context of this work.
Appendix D Executing Quantum Circuits
D.1 Quantum Simulations
D.1.1 Hardware Topology and Swap Network
@@
For the QAOA, each qubit encodes a binary variable of the problem of interest. Therefore we seek the mapping between qubits and variables that will minimize the swap overhead needed to encode the phase separator unitary of Eq. (16). For an arbitrary mapping between $N$ variables and $N$ qubits, an arbitrary QAOA phase separator unitary can be encoded in a swap network of depth $O(N)$ on, for example, a linear or square hardware topology [35, 92].
Efficient Swap Network Finder—
We employ the swap network finder introduced in Ref. [88] and successfully used in Ref. [93]. The hardware topology and phase separator are encoded as graphs $G(V,E)$ and $G^{\prime}(V^{\prime},E^{\prime})$, respectively. Vertices $V^{\prime}$ in the phase separator graph correspond to qubits, and each edge in $E^{\prime}$ corresponds to a term $e^{i\gamma\mathsf{W}_{ij}\hat{Z}_{i}\hat{Z}_{j}/2}\neq\hat{I}$ from Eq. (16), between two qubits $i$ and $j$. The swap network finder is partly about solving a subgraph isomorphism problem. We perform this task using the approach described in App. C.3, based on a Boolean satisfiability (SAT) formulation of the problem. It works as follows: one iteratively updates the connectivity of the hardware topology graph $G$ by applying a set of swap gates. After the swap gates have been applied, one solves a subgraph isomorphism problem on the updated graph. If $G^{\prime}$ can be embedded onto $G$, then one has found a satisfying mapping of variables onto qubits as well as a satisfying set of swap gates, i.e., a satisfying swap network. If $G^{\prime}$ cannot be embedded onto $G$, one applies another set of swap gates to $G$ and repeats the operation until a satisfying swap network is found. It is the responsibility of the user to define the set of swap gates since there is no optimization of the swap network per se. Hence, one starts with no or a small swap network and increases its depth until a satisfying solution is found.
@@
Figure 11: (a) Graph representing a QAOA phase separator made of $N=6$ qubits. (b) Linear chain topology with $6$ qubits and a single layer of swap gates applied on even edges. (c) Mapping of the QAOA phase separator qubits onto the linear chain topology. The extra effective edges introduced by the layer of swap gates are marked in green. A single layer of swap gates enables the embedding of the graph from (a) onto the linear chain topology.
Data—
We map an $N$-variable QAOA phase separator onto an $N$-variable line topology. At each iteration of the swap network builder, we alternatively add a layer of parallel swap gates on even and odd edges of the line topology. This operation is repeated until a satisfying swap network is found. After this step, the quantum circuit encoding the desired QAOA phase separator for a linear qubit topology is fixed. This strategy is applied to an $N=6$-qubit QAOA phase separator in Fig. 11. The swap network builder can find a mapping of the graph nodes onto qubits requiring only a single layer of two-qubit gates. The assignment onto a linear chain is as follows: $[6,1,2,4,5,3]$ with two-qubit operations possible on edges $(6,1)$, $(1,2)$, $(2,4)$, $(4,5)$, and $(5,3)$. From there, a layer of swap gates leads to the following arrangement $[1,6,4,2,3,5]$, enabling two-qubit operations on edges $(1,6)$, $(6,4)$, $(4,2)$, $(2,3)$, $(3,5)$. In particular, edges $(6,4)$ and $(2,3)$ are made available through the swap gates, as shown in Fig. 11, enabling the encoding of the graph onto the linear topology of the qubits. The strategy is guaranteed to succeed with at most $N$ alternating even and odd layers of swap gates [35].
The algorithm for finding an efficient swap network requires a SAT solver for establishing graph isomorphism. The run-times required to find such swap networks are shown in Fig. 10b. Following Eq. (23), the maximum size of a subgraph as a function of the number of QAOA layers $p$ is $1+6(2^{p}-1)$.The distribution of subgraph sizes was discussed in App. C.2.2 and plotted in Fig. 7. The run-times in Fig. 10b were measured on a $64$Gb MacBook Pro with an Apple M1 Max chip at fixed $p$, independent of the size of the subgraphs themselves. We used the SAT solver Glucose [89], based on Minisat [90], through the Python package PySAT [91]. Given that one has to solve a SAT problem, the method becomes prohibitively expensive as $p$ increases. We refer the reader to Ref. [88] for further discussion.
@@
Remarks—
For estimating the run-time of the quantum algorithm in Figs. 2 and 3, we assume that the circuit is mapped onto a hardware-native linear chain of qubits with a default, worst-case-scenario swap network yielding a total number of layers of two-qubit gates proportional to $Np$ for $N$ qubits and $p$ QAOA layers [35]. The quantum algorithm depth (and run-time) does not assume the use of an efficient swap network. In the main text and App. D, a default, worst-case-scenario swap network on a linear chain of qubit is assumed.
D.1.2 Compilation
Figure 12: Construction of algorithmic gates in terms of hardware-native gates on Rigetti’s AnkaaTM-2 superconducting quantum chip. One-qubit gates are defined in Eq. (27). (a) Implementation of $\texttt{Rx}(\phi)$ for arbitrary angles $\phi$. (b) Implementation of the Hadamard gate. (c) Implementation of the two-qubit $\texttt{Rzz}(\phi)$ gate. (d) Implementation of the two-qubit $\texttt{Rzz}(\phi)\times\texttt{SWAP}$ gate.
The native gates on Rigetti’s AnkaaTM-2 are the one-qubit rotation gates about the $x$ axis $\texttt{Rx}(\phi)=\exp(-i\hat{X}\phi/2)$ where $\phi=\pm\pi/2$, and the two-qubit gate $\texttt{ISWAP}=\exp[i\pi(\hat{X}\otimes\hat{X}+\hat{Y}\otimes\hat{Y})/4]$. One-qubit rotations about the $z$ axis, $\texttt{Rz}(\theta\in\mathbb{R})=\exp(-i\hat{Z}\theta/2)$, are implemented virtually. $\hat{X}$, $\hat{Y}$, and $\hat{Z}$ are Pauli operators. Our parameterized circuits have $\texttt{Rx}(\phi)$, $\texttt{Rzz}(\phi)=\exp[-i\phi\hat{Z}\otimes\hat{Z}/2]$, and $\texttt{Rzz}(\phi)\times\texttt{SWAP}$ gates, which are not native to Rigetti’s AnkaaTM-2. We implement these gates using the compilation in Fig. 12 where